# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v3

@@ 6. Case studies
Four taxonomies. The first with default parameters (merge threshold 0.75, minimum category size 2), contains a hierarchy of 4 layers and 23 categories. The second that is more likely to merge (merge threshold 0.5, min category size 2), contains 4 layers and 18 categories. The third that is a larger minimum category size (merge threshold 0.75, min category size 5) contains 3 layers and 13 categories). The fourth that is both more likely to merge and a larger min category size (merge threshold 0.5, min category size 5), contains 3 layers and 9 categories).
6.1. Case study 1: Taxonomy creation
For our first case study, we explore taxonomy creation, the task of organizing a set of items into a category tree to aid sensemaking. This task involves many elements of the accuracy objective, with some degree of creativity: categories need to be correct, but most items can fit into multiple categorizations of different levels of precision. We focus on the ability to control the granularity of these categorizations by controlling the taxonomy shape (Figure 15).
6.1.1. Workflow design
To build Cascade-LLM, our version of Cascade powered by LLMs, we made multiple changes to the original workflow. We highlight some of the most salient modifications. To achieve diverse categorizations, we improved response diversity by giving different prompts to parallel generations and ensuring that every item is input into a category generation step. At first, the categories were too broad. Specifying category qualities via targeted prompts proved unsuccessful, so instead we added a programmatic validation step to delete categories that included “and,” “or,” commas, or slashes. As this hard-coded validation step worked, it was not necessary to include another LLM call. The original workflow gives workers sets of 8 items for which to generate 8 categories. However, the LLM would try categorizing multiple items at once, consistent with prior findings that LLMs struggle to disambiguate multiple subtasks in a prompt (Wu et al., 2023c; Madaan et al., 2023b). To increase diversity, we aligned subtask design to fit model capabilities with one item categorization. Finally, the original workflow also requires parameters for the minimum category size and a merge threshold. To support user guidance through direct manipulation, we dynamically apply these parameters at the last taxonomization stage.
6.1.2. Evaluation setup
@@ 6.1.4. Workflows reduce hallucinated and forgotten items
Cascade-LLM does not forget any items, but 83.72% of zero-shot taxonomies do not include all items. Cascade-LLM cannot hallucinate new items, while zero-shot hallucinates 51 items across the 129 taxonomies. Cascade-LLM taxonomies have similar error rates to zero-shot taxonomies on duplicate categories and missing/incorrect category-item and parent-child pairs, yet are richer taxonomies with more categories and depth (Table 1).
@@ 6.1.5. Workflows have higher costs
We measure generation time and the number of API calls to an LLM. Cascade-LLM takes more time (mean: 2,642.58 sec., std: 655.81) than the zero-shot (mean: 24.35 sec., std: 5.04) and zero-shot-target (mean: 21.69 sec., sd: 4.49) baselines. Crowdsourced Cascade took 43 hours and 3 minutes to taxonomize three item sets (Chilton et al., 2013). Cascade-LLM makes more model calls (mean: 18,898.33, std: 4472.61) than the one call from the zero-shot and zero-shot-target baselines. Most of these calls are short in terms of token length, as they compare one item to one category. Crowdsourced Cascade reports 1760 calls to workers per iteration. For our goals, the increased cost in time and computation enables direct manipulation on more detailed categorizations without item forgetting or hallucination.
@@ 6.2. Case study 2: Text shortening
The next case study is the task of shortening text. Unlike summarization, shortening the text involves making edits to create a condensed version of the text that attempts to preserve the full meaning. Shortening is a task on the creative side, but requiring elements of accuracy. Shortened versions must stay faithful to the original text, and there are many options of what phrases to shorten and how to shorten them. By implementing a workflow to generate many shortening options, an end user can control the final length of a paragraph dynamically.
@@ 6.2.1. Workflow design
To shorten text we adapt the Find-Fix-Verify workflow (Figure 16) introduced in Soylent (Bernstein et al., 2010). In the “Find” step, crowdworkers select areas that might be shortened. If at least two workers agree on a text span, that span progresses to the “Fix” step. In the “Fix” step, crowdworkers make edits to shorten the text. In the final “Verify” step, more crowdworkers verify the possible changes by rejecting by majority vote any edits deemed grammatically incorrect or that change the meaning of the sentence. These steps respectively map to the focus, transform, and evaluate subtask types in our design space. The workflow outputs edit options for multiple spans in the paragraph, which in turn may have multiple possible edits. Soylent defaults to showing the shortest possible option, but it also provides a length slider for the user to adjust the final output as desired.
We made multiple adjustments to produce our Soylent-LLM workflow. First, we ran into context limits and parallelization constraints with longer texts. In response, we apply a dynamic approach that splits the workflow into chunks of 10 sentences and applies the workflow to each. Second, Soylent allows workers to fix inputs by editing, merging, and deleting phrases. LLMs performed better when we split these tasks into separate prompts, so we further decompose Find and Fix into finer grained editing, merging, and deleting steps to get an align subtask size to the model’s capabilities. Finally, in the Find step the LLM would sometimes “hallucinate” phrases to shorten that do not exist in the text or would change quotes. To mitigate these errors, we add several hard-coded validation steps: we ensure that the phrase to replace exists in the input text, that the replacement is shorter, and that no quotes were altered or added.
Recent work also implemented three unique variations of a Find-Fix-Verify workflow using LLM chains (Wu et al., 2023c). Their evaluation, using student preferences, found mixed results on effectiveness for creating better shortenings than a zero-shot baseline. We evaluate our workflow along different dimensions: controllability of output length, fluency, grammaticality, what type of content is deleted, and workflow resource usage. Unlike this prior work, we also retain the parallel generation and validation steps from the original paper.
@@ 6.2.2. Evaluation setup
Two graphs side by side. The first graph is a line graph graphing the target number of words on the X axis from 100-500 and percent error on the y axis from -75% error to 250% error. The graph contains 4 lines. On short texts, the zero-shot-target gets the lowest error followed by zero-shot, Soylent-LLM and zero-shot-ffv. On longer texts Soylent-LLM has approximately 0% error. Zero-shot has the highest error, and zero-shot-target and zero-shot-ffv have approximately the same error at 50%/ The 95% confidence intervals are narrow. The second graph is a scatter plot with the X axis the target percent length of original ranging from 0 to 100 and with the Y axis the percentage error from -100 to 500%. There are 4 colors of points. The highest percent errors are for small target lengths across all 4 baselines. As the target percentage increases, Soylent-LLM converges to close to zero percent error, while the three zero-shot baselines trend negative.
@@ 6.2.4. Workflow outputs have abrupt transitions but lower perplexity
We measure the average perplexity of output texts as a proxy for fluency. We use a different language model (Facebook’s OPT-1.3b) than the worker model (OpenAI’s gpt-3.5-turbo) for an external evaluation. We also perform a qualitative analysis of the shortening procedure. We use the five texts used for evaluation in the original Soylent paper (Bernstein et al., 2010). We compare the edits made with the shortest output from Soylent-LLM, the baseline conditions, and the crowdsourced condition’s output provided by the authors of the original paper. Without knowing the condition, we coded the outputs for fluency and grammatical errors, following the qualitative evaluation procedure of the original paper. We additionally coded what content changed during shortening.
@@
As with the original crowdsourcing workflow, some shortenings diminished fluency or introduced grammatical errors. Zero-shot, zero-shot-ffv and Soylent-LLM shortenings all included passive voice, but were otherwise grammatical. However, Soylent-LLM and zero-shot-ffv both could contain abrupt endings or transitions. This abruptness is due to separately editing parts of the input text. Crowdsourced Soylent induced grammatical errors as well as abrupt transitions. We found that zero-shot and Soylent-LLM both shorten by removing unnecessary phrasing and deleting examples and justifications. Zero-shot more successfully merges sentences than Soylent-LLM, while Soylent-LLM changes less text and more successfully retains the user’s voice, rather than formalizing it. Soylent-LLM’s changes affect a larger portion of the text than Crowdsourced Soylent. In sum, Crowdsourced Soylent made 42 edits to phrases and 12 to entire sentences, while Soylent-LLM made 12 edits to phrases and 36 to entire sentences.
@@ 6.2.5. Workflows have higher costs than zero-shot prompting
Soylent-LLM takes much longer (mean: 105.44 sec., std: 60.16) than the zero-shot (mean: 9.09 sec., std: 2.92), zero-shot-target (mean: 13.65 sec., std: 5.47), and zero-shot-ffv (mean: 43.55 sec., std: 23.16) baselines. Crowdsourced Soylent takes longest with a reported range of 2,640–29,340 sec. (44-489 min.) (Bernstein et al., 2010). Compared to the one call to the model for the zero-shot baselines, one Soylent-LLM run averages 127.11 calls to the model (std. 63.56). Crowdsourced Soylent reports incorporating 158–362 workers.
@@ 6.3. Case study 3: Short story generation
The final case study concerns story generation given a short text prompt. This task lies on the open-ended creativity side of the spectrum, as short stories have a wide design space of possible plots, characters, messages, imagery, and lessons. In this task, user guidance of the story involves higher level control of different plot elements, with the ability to see counterfactuals (e.g., if a story element is included or not). We implement a workflow that, compared to zero-shot approaches, enables this exploration more precisely, while also creating stories that are preferred by crowdsourced raters for imagery, originality, and style.
@@ 6.3.1. Workflow design
We adapt the Mechanical Novel (MN) (Kim et al., 2017) workflow for this task. This workflow involves three main steps (Figure 18). The first generate subtask “initializes” a story with multiple scenes based on a given story prompt (e.g., “Malcolm finds himself alone in a runaway hot air balloon and accidentally travels to a city in the sky.”). Then, the “reflect” and “revise” steps are iterated five times in a loop. In the “reflect” step, workers consider what changes might helpfully adjust the story. In the “revise” step, workers implement those changes. Each step has subtasks for parallel generation of suggestions and edits, as well as evaluate subtasks voting for the best generation.
We adapted the Crowdsourced MN to MN-LLM by incorporating fine-grained user guidance with manipulatable outputs. At the conclusion of the workflow, instead of receiving just one or more separate short stories, the user can interact with different combinations of edits to compare stories with different quality combinations and compare counterfactuals on editing decisions. To enable these outputs, we run the “revise” stage on texts of all combinations of suggestions from the “reflect” stage.
Beyond introducing control of story elements, the largest challenge was managing token context limitations. Getting response diversity with iterative improvement on texts often increases the text length (Little et al., 2010a). To account for context length, we generate stories with four scenes rather than six and iterate through scenes rather than informing workers to select scenes. In our design space this change maps to using a redundant architecture rather than a sequential architecture with a focus subtask. At the beginning of the “revise” step, we add a generation subtask that summarizes the current story to provide abbreviated context in the worker instructions. We included a hard-coded validation step with an evaluate subtask to disqualify edits that were more than 1.5 times the input text length. Unfortunately, this disqualified many highly descriptive texts and sometimes meant that no edits were made because all suggestions were too long.
| MN-LLM | 124 | 97 | 93 | 137 | 118 | 97 | 112 |
| Zero-shot | 76 | 103 | 107 | 63 | 82 | 103 | 88 |
@@ 6.3.5. The MN-LLM workflow improves imagery, originality, and style
Following the evaluation methods of the MN paper, for each story prompt we gave the output from MN-LLM and the zero-shot baseline, in randomized order, to 200 workers on Amazon Mechanical Turk using EasyTurk (Krishna, 2019). We asked crowdworkers to vote between stories from the Soylent-LLM and zero-shot baselines along seven story writing dimensions (Kim et al., 2017). We paid workers $1.50 USD, the equivalent of $18 per hour.
@@ 6.3.7. Workflows have higher costs
We measure the time and number of API calls required to run the workflow compared to running the baselines. Compared to zero-shot (mean: 21.88 sec., std: 5.89), MN-LLM takes much longer in seconds to complete (mean: 1,452.97 sec., std: 930.26). Compared to zero-shot’s one call, MN-LLM makes 2,398 LLM calls on average. The crowdsourced Mechanical Novel paper does not report latency or number of workers (Kim et al., 2017).
@@ 7. Discussion
Thus, strategies and their effect on objectives transfer directly from crowdsourcing to LLMs. However, the tactic designs to support strategies can vary significantly, and could benefit from improved development and testing processes. Understanding how to effectively use tactics to implement strategies is not only a more focused goal than the connection between tactics and objectives, but it is also more generalizeable. For example, learning how to effectively get LLMs to validate their own outputs would be applicable to many different tasks and objectives.
@@ 7.1. LLM stochasticity is different than human diversity
One of the key benefits of crowdsourcing workflows is that they incorporate multiple people, such that crowdworkers can fill in each other’s deficits (Huang et al., 2017; Nebeling et al., 2016; Kim and Monroy-Hernandez, 2016; Zhang et al., 2012). Therefore, requesting multiple crowdworkers for the same or sequential subtasks inherently leads to differing outputs and considerations. The diversity of perspectives happens inherently by querying different workers but is limited to available crowdworkers and thus affected by demographic and geographic disparities on crowdsourcing platforms (Difallah et al., 2018).
@@
LLMs also have the capability to output different responses, as their training data includes information produced by many people on the Internet (Bender et al., 2021). To induce diversity, prior work has tried methods like adjusting prompts, temperature, and model types (Parameswaran et al., 2023; Park et al., 2022; Du et al., 2023; Arora et al., 2022). However, ongoing work has found that there are elements of diversity in people that are not evident in LLMs. For example, LLMs struggle to output pluralistic responses (Sorensen et al., 2024), are limited to the outputs of people with access to the Internet (Bender et al., 2021), and cannot represent identity groups (Wang et al., 2024). Our case studies suggest that LLMs need more variation in prompting and explicit intervention to get diverse responses, although we did not test this systematically. Intuitively, this makes sense because there is great diversity in the human population, but models are known to mimic the majority voice to minimize their objectives, ignoring marginalized opinions that make up a smaller percentage of the training data (Gordon et al., 2022).
@@
Using a high temperature and prompting the LLM with a variety of different instructions can improve diversity somewhat but is unlikely to match that of crowds. Although this difference between workers may at some level seem obvious, it impacts the tactics needed to effectively use diverse responses. Diverse responses are used both to source variety and to check agreement. In crowdsourcing, the underlying tactic behind both of these goals is identical: run the same subtask with multiple workers. However, different tactics can better explicitly address these applications for LLMs.
@@
Variety. Diverse responses are often used to support creative tasks with a variety of outputs. Consider one step in Mechanical Novel that generates suggestions for updating the story. As this is an iterative workflow, we want a variety of story update suggestions both at each iteration and across iterations. Unfortunately, even with variety in generation prompt wording, we found a low diversity of suggestion content (Table 4 “Zero-shot” column). After repeated challenges in prompt engineering to increase the variety of suggestions, we eventually intervened with an added validation step to ensure diversity of story updates. This approach was effective, but inefficient, as it filters out many generated responses. Role-based communicative architectures are also promising in giving prompts extra power to increase variety. Although giving identity based roles should treated with caution (Wang et al., 2024), giving other roles to LLMs could be a more powerful way of inducing a variety of responses (Park et al., 2022). Many roles in crowdsourcing workflows were based on organizational positions for relevant domains such as journalism (Agapie et al., 2015) or project development (Valentine et al., 2017).
@@
To source variety, the best choice of redundant architecture can differ when using crowdworkers or LLMs. Prior crowdsourcing work has found that parallel and iterative architectures source different degrees of diversity and quality of responses in crowdsourcing depending on the task (Goto et al., 2016; Little et al., 2010a; Kittur et al., 2011; André et al., 2014). For example, iterative architectures have less variation and a higher average quality of outputs, but have a lower maximum quality output (Little et al., 2010a). Some of these tradeoffs will hold with LLMs, but differences may emerge as well. For instance, some work suggests that, relative to crowdworkers who experience anchoring bias on iterative architectures, LLMs can perform divergent thinking to generate something explicitly different than provided examples (Wu et al., 2023c, a). Similar studies on LLM chains would better inform the effect of different redundant architectures on variety and quality.
@@
Agreement. Crowdsourcing work has used agreement among crowdworkers as a proxy for confidence vs uncertainty (Liem et al., 2011; Chang et al., 2017). This agreement can be used to validate the quality of outputs, so understanding how agreement relates to quality is of critical importance. Here, we focus on agreement among outputs of identical tasks, but the difference between a different person validating the outputs of another, and a model sequentially validating its own outputs, should also be considered.

### v4

@@ 7.1.4. Workflows reduce hallucinated and forgotten items
Cascade-LLM does not forget any items, but 83.72% of zero-shot taxonomies do not include all items. Cascade-LLM cannot hallucinate new items, while zero-shot baselines hallucinate 51 items across the 129 taxonomies. Cascade-LLM taxonomies have similar error rates to zero-shot baselines’ taxonomies on duplicate categories and missing/incorrect category-item and parent-child pairs, yet are richer taxonomies with more categories and depth (Table 3).
@@ 7.1.5. Workflows have higher costs
We measure generation time and the number of API calls to an LLM. Cascade-LLM takes more time (mean: 2,642.58 sec., std: 655.81) than the zero-shot (mean: 24.35 sec., std: 5.04) and zero-shot-target (mean: 21.69 sec., sd: 4.49) baselines. Crowdsourced Cascade took 43 hours and 3 minutes to taxonomize three item sets (Chilton et al., 2013). Cascade-LLM makes more model calls (mean: 18,898.33, std: 4472.61) than the one call from the zero-shot and zero-shot-target baselines. Most of these calls are short in terms of token length, as they compare one item to one category. Crowdsourced Cascade reports 1760 calls to crowdworkers per iteration. For our goals, the increased cost in time and computation enables direct manipulation of more detailed categorizations without item forgetting or hallucination.
@@ 7.2. Case study 2: Text shortening
The next case study is the task of shortening text. Unlike summarization, shortening the text involves making edits to create a condensed version of the text that attempts to preserve the full meaning. Shortening is a task primarily on the subjectivity side, but requiring elements of objectivity. Shortened versions must stay faithful to the original text, and there are many options for what phrases to shorten and how to shorten them. By implementing a workflow to generate many shortening options, an end user can control the final length of a paragraph dynamically.
@@ 7.2.1. Workflow design
To shorten text we adapt the Find-Fix-Verify workflow (Figure 18) introduced in Soylent (Bernstein et al., 2010). In the “Find” step, crowdworkers select areas that might be shortened. If at least two actors agree on a text span, that span progresses to the “Fix” step. In the “Fix” step, crowdworkers make edits to shorten the text. In the final “Verify” step, more crowdworkers verify the possible changes by rejecting by majority vote any edits deemed grammatically incorrect or that change the meaning of the sentence. These steps respectively map to the focus, improve, and evaluate subtask types in our design space. The workflow outputs edit options for multiple spans in the paragraph, which when combined create multiple possible outcome texts. Soylent defaults to showing the shortest possible option, but it also provides a length slider for the user to adjust the final output as desired. The slider enables the user to test different shortening phrasing options and lengths through a simple interaction.
We made multiple adjustments to produce our Soylent-LLM workflow. First, we ran into context limits and parallelization constraints with longer texts. In response, we apply a dynamic approach that splits the workflow into chunks of 10 sentences and applies the workflow to each. Second, Soylent allows actors to fix inputs by editing, merging, and deleting phrases. LLMs performed better when we split these tasks into separate prompts, so we further decompose Find and Fix into finer grained editing, merging, and deleting steps to get an aligned subtask size to the model’s capabilities. Finally, in the Find step, the LLM would sometimes “hallucinate” phrases to shorten that do not exist in the text or would change quotes. To mitigate these errors, we add several hard-coded validation steps: we ensure that the phrase to replace exists in the input text, that the replacement is shorter, and that no quotes were altered or added.
Recent work also implemented three unique variations of a Find-Fix-Verify workflow using LLM chains (Wu et al., 2023b). Their evaluation, using student preferences, found mixed results on effectiveness for creating better shortenings than a zero-shot baseline. We evaluate our workflow along different dimensions: controllability of output length, fluency, grammaticality, what type of content is deleted, and workflow resource usage. Unlike this prior work, we also retain the parallel generation and validation steps from the original paper.
@@ 7.2.2. Evaluation setup
Two graphs side by side. The first graph is a line graph graphing the target number of words on the X axis from 100-500 and percent error on the y axis from -75% error to 250% error. The graph contains 4 lines. On short texts, the zero-shot-target gets the lowest error followed by zero-shot, Soylent-LLM, and zero-shot-ffv. On longer texts, Soylent-LLM has approximately 0% error. Zero-shot has the highest error, and zero-shot-target and zero-shot-ffv have approximately the same error at 50%/ The 95% confidence intervals are narrow. The second graph is a scatter plot with the X axis the target percent length of the original ranging from 0 to 100 and with the Y axis the percentage error from -100 to 500%. There are 4 colors of points. The highest percent errors are for small target lengths across all 4 baselines. As the target percentage increases, Soylent-LLM converges to close to zero percent error, while the three zero-shot baselines trend negative.
@@ 7.2.4. Workflow outputs have abrupt transitions but lower perplexity
We measure the average perplexity of output texts as a proxy for fluency. We use a different language model (Facebook’s OPT-1.3b) than the actor model (OpenAI’s gpt-3.5-turbo) for an external evaluation. We also perform a qualitative analysis of the shortening procedure. We use the five texts used for evaluation in the original Soylent paper (Bernstein et al., 2010). We compare the edits made with the shortest output from Soylent-LLM, the baseline conditions, and the crowdsourced condition’s output provided by the authors of the original paper. Without knowing the condition, we coded the outputs for fluency and grammatical errors, following the qualitative evaluation procedure of the original paper. We additionally coded what content changed during shortening.
@@
As with the original crowdsourcing workflow, some shortenings diminished fluency or introduced grammatical errors. Zero-shot, zero-shot-ffv, and Soylent-LLM shortenings all included passive voice but were otherwise grammatical. However, Soylent-LLM and zero-shot-ffv both could contain abrupt endings or transitions. This abruptness is due to separately editing parts of the input text. Crowdsourced Soylent induced grammatical errors as well as abrupt transitions. We found that zero-shot and Soylent-LLM both shorten by removing unnecessary phrasing and deleting examples and justifications. Zero-shot more successfully merges sentences than Soylent-LLM, while Soylent-LLM changes less text and more successfully retains the user’s voice, rather than formalizing it. Soylent-LLM’s changes affect a larger portion of the text than Crowdsourced Soylent. In sum, Crowdsourced Soylent made 42 edits to phrases and 12 to entire sentences, while Soylent-LLM made 12 edits to phrases and 36 to entire sentences.
@@ 7.2.5. Workflows have higher costs than zero-shot prompting
Soylent-LLM takes much longer (mean: 105.44 sec., std: 60.16) than the zero-shot (mean: 9.09 sec., std: 2.92), zero-shot-target (mean: 13.65 sec., std: 5.47), and zero-shot-ffv (mean: 43.55 sec., std: 23.16) baselines. Crowdsourced Soylent takes the longest with a reported range of 2,640–29,340 sec. (44-489 min.) (Bernstein et al., 2010). Compared to the one call to the model for the zero-shot baselines, one Soylent-LLM run averages 127.11 calls to the model (std. 63.56). Crowdsourced Soylent reports incorporating 158–362 crowdworkers.
7.2.6. Chaining demonstrates tradeoffs with Chain of Thought
Zero-shot-ffv is a Chain of Thought (CoT) (Wei et al., 2022) baseline, in which the LLM is prompted to use the steps of the Find-Fix-Verify chain all in one prompt-response set. As previously discussed, zero-shot-ffv follows the quality trends of Soylent-LLM more closely than those of the zero-shot baseline. Compared to Soylent-LLM, the zero-shot-ffv baseline’s quality and cost are lower. Therefore, the relative utility of doing a chain or a CoT depends on the restrictions on cost and the importance of high quality. We did not do a CoT baseline for the other two case studies, as context lengths did not enable the chain to be represented in one prompt and response. Therefore, another benefit of chains over CoT is the capability to do tasks and follow logical flows outside the context limit of an LLM. These restrictions will change as the context limits of LLMs grow.
@@ 7.3. Case study 3: Short story generation
The final case study concerns story generation given a short text prompt. This task lies on the open-ended subjectivity side of the spectrum, as short stories have a wide design space of possible plots, characters, messages, imagery, and lessons. In this task, user guidance of the story involves higher level control of different plot elements, with the ability to see counterfactuals (e.g., if a story element is included or not). We implement a workflow that, compared to zero-shot approaches, enables this exploration more precisely, while also creating stories that are preferred by crowdsourced raters for imagery, originality, and style.
@@ 7.3.1. Workflow design
We adapt the Mechanical Novel (MN) (Kim et al., 2017) workflow for this task. This workflow involves three main steps (Figure 20). The first generate subtask “initializes” a story with multiple scenes based on a given story prompt (e.g., “Malcolm finds himself alone in a runaway hot air balloon and accidentally travels to a city in the sky.”). Then, the “reflect” and “revise” steps are iterated five times in a loop. In the “reflect” step, actors consider what changes might helpfully adjust the story. In the “revise” step, actors implement those changes. Each step has subtasks for parallel generation of suggestions and edits, as well as evaluate subtasks voting for the best generation. Crowdworkers additionally focus editing attention on the most important scenes.
We adapted the Crowdsourced MN to MN-LLM by incorporating fine-grained user guidance with manipulatable outputs. At the conclusion of the workflow, instead of receiving just one or more separate short stories, the user can mark check-boxes on an interface selecting different combinations of edits to view stories with different qualities and compare counterfactuals on editing decisions. To enable these outputs, we run the “revise” stage on texts of all combinations of suggestions from the “reflect” stage.
Beyond introducing control of story elements, the largest challenge was managing token context limitations. Getting response diversity with iterative improvement on texts often increases the text length (Little et al., 2010a). To account for context length, we generate stories with four scenes rather than six and iterate through scenes rather than informing actors to select scenes. In our design space this change maps to using a redundant iterative architecture rather than a sequential architecture with a focus subtask. At the beginning of the “revise” step, we add a generation subtask that summarizes the current story to provide abbreviated context in the worker instructions. We included a hard-coded validation step with an evaluate subtask to disqualify edits that were more than 1.5 times the input text length. Unfortunately, this disqualified many highly descriptive texts and sometimes meant that no edits were made because all suggestions were too long.
@@ 7.3.5. The MN-LLM workflow improves imagery, originality, and style
Following the evaluation methods of the MN paper, for each story prompt we gave the output from MN-LLM and the zero-shot baseline, in randomized order, to 200 crowdworkers on Amazon Mechanical Turk using EasyTurk (Krishna, 2019). We asked crowdworkers to vote between stories from the Soylent-LLM and zero-shot baselines along seven story writing dimensions (Kim et al., 2017). We paid crowdworkers $1.50 USD, the equivalent of $18 per hour.
@@ 7.3.7. Workflows have higher costs
We measure the time and number of API calls required to run the workflow compared to running the baselines. Compared to zero-shot (mean: 21.88 sec., std: 5.89), MN-LLM takes much longer in seconds to complete (mean: 1,452.97 sec., std: 930.26). Compared to zero-shot’s one call, MN-LLM makes 2,398 LLM calls on average. The crowdsourced Mechanical Novel paper does not report latency or the number of crowdworkers (Kim et al., 2017).
@@ 8. Discussion
Thus, strategies and their effect on objectives transfer directly from crowdsourcing to LLMs. However, the tactic designs to support strategies can vary significantly and could benefit from improved development and testing processes. Understanding how to effectively use tactics to implement strategies is not only a more focused goal than the connection between tactics and objectives, but it is also more generalizable. For example, learning how to effectively get LLMs to validate their own outputs would be applicable to many different tasks and objectives.
@@ 8.1. LLM stochasticity is different than human diversity
One of the key benefits of crowdsourcing workflows is that they incorporate multiple people, such that crowdworkers can fill in each other’s deficits (Huang et al., 2017; Nebeling et al., 2016; Kim and Monroy-Hernandez, 2016; Zhang et al., 2012). Therefore, requesting multiple crowdworkers for the same or sequential subtasks inherently leads to differing outputs and considerations. The diversity of perspectives happens inherently by querying different actors but is limited to available crowdworkers and thus affected by demographic and geographic disparities on crowdsourcing platforms (Difallah et al., 2018).
@@
LLMs also have the capability to output different responses, as their training data includes information produced by many people on the internet (Bender et al., 2021). To induce diversity, prior work has tried methods like adjusting prompts, temperature, and model types (Parameswaran et al., 2023; Park et al., 2022; Du et al., 2023; Arora et al., 2023). In some cases, one can automatically calculate the optimal mix of LLMs for a task (Madaan et al., 2023). Our case studies suggest that LLMs need more variation in prompting and explicit intervention to get diverse responses, although we did not test this systematically. Intuitively, this makes sense because there is great diversity in the human population, but models are known to mimic the majority voice to minimize their objectives, ignoring marginalized opinions that are a smaller percentage of the training data (Gordon et al., 2022).
@@
Using a high temperature and prompting the LLM with a variety of different instructions can improve diversity somewhat but is unlikely to match that of crowds. Although this difference between actors may at some level seem obvious, it impacts the tactics needed to effectively use diverse responses. Diverse responses are used both to source variety and to check agreement. In crowdsourcing, the underlying tactic behind both of these goals is identical: run the same subtask with multiple actors. However, different tactics can better explicitly address these applications for LLMs.
@@
Variety. Diverse responses are often used to support tasks with a variety of subjective outputs. Consider one step in Mechanical Novel that generates suggestions for updating the story. As this is an iterative workflow, we want a variety of story update suggestions both at each iteration and across iterations. Unfortunately, even with variety in generation prompt wording, we found a low diversity of suggestion content (Table 5 “Zero-shot” column). After repeated challenges in prompt engineering to increase the variety of suggestions, we eventually intervened with an added validation step of multiple evaluate subtasks to ensure diversity of story updates. This approach was effective, but inefficient, as it filters out many generated responses. Role-based communicative architectures are also promising in giving prompts extra power to increase variety. Although giving identity based roles should treated with caution (Wang et al., 2024), giving other roles to LLMs could help induce a variety of responses (Park et al., 2022). Many roles in crowdsourcing workflows were based on organizational positions for relevant domains such as journalism (Agapie et al., 2015) or project development (Valentine et al., 2017).
@@
To source variety, the best choice of redundant architecture can differ when using crowdworkers or LLMs. Prior crowdsourcing work has found that parallel and iterative architectures source different degrees of diversity and quality of responses in crowdsourcing depending on the task (Goto et al., 2016; Little et al., 2010a; Kittur et al., 2011; André et al., 2014). For example, iterative architectures have less variation and a higher average quality of outputs but have a lower maximum quality output (Little et al., 2010a). Some of these tradeoffs will hold with LLMs, but differences may emerge as well. For instance, some work suggests that, relative to crowdworkers who experience anchoring bias on iterative architectures, LLMs can perform divergent thinking to generate something explicitly different than provided examples (Wu et al., 2023b, a). Similar studies on LLM chains would better inform the effect of different redundant architectures on variety and quality.
@@
Agreement. Crowdsourcing work has used agreement among crowdworkers as a proxy for confidence vs uncertainty (Liem et al., 2011; Chang et al., 2017). This agreement can be used to validate the quality of outputs, so understanding how agreement relates to quality is of critical importance. Here, we focus on agreement among outputs of identical tasks, but the difference between a different person validating the outputs of another and a model sequentially validating its own outputs should also be considered.